### v1

# A Benchmark of Causal vs. Correlation AI for Predictive Maintenance

Abstract.
Predictive maintenance in manufacturing environments presents a challenging optimization problem characterized by extreme cost asymmetry, where missed failures incur costs roughly fifty times higher than false alarms. Conventional machine learning approaches typically optimize statistical accuracy metrics that do not reflect this operational reality and cannot reliably distinguish causal relationships from spurious correlations. This study evaluates eight predictive models, ranging from baseline statistical approaches to formal causal inference methods, on a dataset of 10,000 CNC machines with a 3.3 percent failure prevalence. The formal causal inference model (L5) achieved estimated annual cost savings of 1.16 million USD (a 70.2 percent reduction), outperforming the best correlation-based decision tree model (L3) by approximately 80,000 USD per year. The causal model matched the highest observed recall (87.9 percent) while reducing false alarms by 97 percent (from 165 to 5) and attained a precision of 92.1 percent, with a train-test performance gap of only 2.6 percentage points. These results indicate that causal AI methods, when combined with domain knowledge, can yield superior financial outcomes and more interpretable predictions compared to correlation-based approaches in predictive maintenance applications.
[…]
1. Contributions
This study makes the following contributions to predictive maintenance optimization:
  • Systematic comparison of causal AI vs. correlation-based machine learning: We evaluate eight models ranging from basic baselines to advanced causal methods on a dataset of 10,000 CNC machines from the UCI Machine Learning Repository [1], providing the first comprehensive benchmark of formal causal inference against traditional ML approaches for predictive maintenance.
  • Business-aligned optimization framework: We demonstrate that models optimized explicitly for business costs (accounting for the 50:1 cost asymmetry between false negatives and false positives) substantially outperform accuracy-optimized approaches, highlighting the critical importance of cost-sensitive learning in industrial applications.
  • Superior financial performance through causal modeling: The formal causal inference model (Model L5) achieved 1,157,500 USD in annual cost savings (70.2% reduction vs. reactive maintenance), outperforming the best correlation-based model (Model L3) by 80,000 USD annually while maintaining identical recall (87.9%).
  • Dramatic reduction in false alarms: The causal model reduced false positives by 97% (from 165 to 5) compared to the best correlation-based approach, achieving 92.1% precision and minimizing operational disruption while maintaining safety through high recall.
  • Enhanced interpretability and actionable insights: Unlike black-box correlation models, the causal approach provides root-cause explanations for each prediction (thermal stress, power overload, or mechanical overstrain), enabling targeted maintenance interventions rather than generic inspections.
  • Robust generalization under distribution shift: Causally-derived features demonstrated stronger generalization (2.6 percentage point train-test gap) compared to raw sensor correlations (4.1 percentage points), consistent with theoretical expectations that causal relationships are invariant under operational changes.
  • Analysis of Bayesian ensemble methods: The PyMC BART model (L7) validates that automatic feature generation can approach causal performance, but without graph-informed constraints it issues substantially more false positives, demonstrating that Bayesian ensembles still benefit from structural causal knowledge.
  • Validation of causal AI value proposition: Results demonstrate that formal causal inference that relies on domain expertise to construct causal graphs delivers superior financial performance, operational efficiency, and interpretability compared to state-of-the-art correlation-based approaches in predictive maintenance applications.
2. Introduction
Manufacturing faces a maintenance dilemma. Reactive maintenance leads to costly downtime, while overly cautious preventive maintenance wastes resources on healthy equipment. The core challenge is an extreme cost asymmetry across industrial equipment: a missed failure can cost tens of thousands of dollars (unplanned downtime, emergency repairs, production losses), while a false alarm costs only hundreds of dollars (scheduled inspection, standard parts). For the CNC machines analyzed in this study, we use representative costs of 25,000 USD for missed failures and 500 USD for false alarms, reflecting typical manufacturing industry values where the cost ratio between unplanned downtime and preventive maintenance ranges from 20:1 to 100:1 depending on equipment criticality. Table 1 summarizes the four possible prediction outcomes, their associated costs, and business impact.
Outcome Description Cost Business Impact
True Positive (TP) Correctly predict failure 5,000 USD Scheduled repair prevents downtime
False Positive (FP) Wrongly predict failure 500 USD Wasted inspection/parts
False Negative (FN) Miss actual failure 25,000 USD Emergency downtime + repairs
True Negative (TN) Correctly predict no failure 0 USD Normal production
Table 1. Prediction outcomes, costs, and business impact. Cost estimates reflect typical manufacturing industry values where unplanned downtime incurs significant production losses, emergency labor, and expedited parts costs, while preventive maintenance requires only scheduled labor and standard parts procurement [8].
Standard machine learning models optimize for statistical accuracy metrics such as accuracy, precision, recall, F1-score, and AUC-ROC, which either ignore class imbalance or treat all prediction errors equally. Accuracy is misleading in imbalanced datasets—a model that always predicts “no failure” achieves 96.7% accuracy on our data while providing zero business value. Precision focuses on minimizing false alarms but may miss critical failures, while recall prioritizes catching all failures but may trigger hundreds of unnecessary inspections. F1-score balances precision and recall through their harmonic mean, and AUC-ROC evaluates the model’s ranking ability across all possible thresholds, but neither metric accounts for the actual business costs of different error types. All these metrics fail to capture the 50:1 cost ratio between false negatives and false positives that is typical for manufacturing equipment, where unplanned downtime costs far exceed scheduled maintenance [8]. Furthermore, correlation-based models identify patterns that co-occur with failures but cannot distinguish causation from spurious relationships, leading to poor generalizability and a lack of actionable insights.
Our goal is to determine if Causal AI can deliver superior business value over state-of-the-art, cost-optimized correlation-based ML in a real-world predictive maintenance scenario.
3. Dataset and Experimental Design
Our analysis utilizes the UCI Machine Learning Repository’s CNC Machine Predictive Maintenance dataset, which contains operational data from 10,000 machines over their production lifecycles. Each machine record includes six sensor measurements collected during normal operation as described in Table 2:
Feature Description
Air temperature Measured in Kelvin
Process temperature Measured in Kelvin
Rotational speed Revolutions per minute
Torque Newton-meters
Cumulative tool wear Minutes of operation
Machine type classification Low, Medium, or High specification
Table 2. Feature descriptions
A snippet of the data is shown in Table 3.
UDI Product ID Type Air temp [K] Process temp [K] Rot. speed [rpm] Torque [Nm] Tool wear [min] Machine failure TWF HDF PWF OSF RNF
1 M14860 M 298.1 308.6 1551 42.8 0 0 0 0 0 0 0
2 L47181 L 298.2 308.7 1408 46.3 3 0 0 0 0 0 0
3 L47182 L 298.1 308.5 1498 49.4 5 0 0 0 0 0 0
4 L47183 L 298.2 308.6 1433 39.5 7 0 0 0 0 0 0
5 L47184 L 298.2 308.7 1408 40.0 9 0 0 0 0 0 0
Table 3. Dataset snippet showing the first five rows of the CNC machine predictive maintenance dataset.
The dataset documents five distinct failure modes that can occur in CNC machines, as summarized in Table 4:
Symbol Failure Type Description Instances
TWF Tool Wear Failure Tool replacement or failure at randomly selected wear time 120
HDF Heat Dissipation Failure Process failure due to temperature differential (process minus air) 115
PWF Power Failure Process failure due to mechanical power (torque $\times$ rotational speed in rad/s) 95
OSF Overstrain Failure Process failure due to tool wear $\times$ torque exceeds threshold 98
RNF Random Failure 0.1% probability of failure regardless of process parameters (unpredictable) 5
Table 4. Failure types with symbols, descriptions, and instance counts. Definitions follow the dataset documentation [1].
3.1. Target Definition and Rationale
For this analysis, we focus exclusively on deterministic failures—those failure modes that can theoretically be predicted from sensor measurements and operating conditions. Specifically, our prediction target combines Tool Wear Failure, Heat Dissipation Failure, Power Failure, and Overstrain Failure, while explicitly excluding Random Failure events. Random failures are defined in the dataset as having a fixed 0.1% probability of occurrence regardless of process parameters—they are statistically independent of all sensor readings and represent truly unpredictable component failures (only 5 instances in 10,000 data points). This exclusion reflects a fundamental principle: predictive models should be evaluated on phenomena they can reasonably be expected to predict. Including random failures would unfairly penalize model performance on events that contain no predictive signal by design.
After applying this filter, our dataset contains 9,995 total observations. We employ a stratified 80-20 train-test split, preserving the class distribution in both sets. The test set comprises 2,000 machines, of which 66 experience deterministic failures—a failure rate of 3.3%. This severe class imbalance poses significant challenges for standard machine learning approaches and underscores why accuracy-based metrics are inappropriate for evaluation.
  • Size: 10,000 machines with sensor readings
  • Features: Air temperature, process temperature, rotational speed, torque, tool wear, machine type (L/M/H)
  • Target: Deterministic failures excluding random failures (TWF + HDF + PWF + OSF)
  • Test Set: 2,000 machines (20% stratified split) with 66 failures (3.3% prevalence)
  • Baseline Cost: 1,650,000 USD (reactive maintenance: 66 failures × 25,000 USD)
3.2. Methodology
3.2.1. Success Metrics
The metrics we consider are the following:
  1. 1.
    Total annual cost: The business objective is to minimize the total cost of ownership of the machine
  2. 2.
    Savings: Dollar reduction vs. reactive maintenance baseline
  3. 3.
    Recall: Percentage of actual failures caught (which can be critical for safety)
  4. 4.
    Precision: Percentage of alarms that are real failures (which can be need to minimize waste)
  5. 5.
    Generalization: Train vs. test performance gap
3.2.2. Correlation-Based Models (L0-L3)
We developed four correlation-based models of increasing sophistication, representing the spectrum of traditional machine learning techniques applied to predictive maintenance.
Model L0: No-Skill Baseline. The no-skill model serves as the minimal performance floor. This model predicts “no failure” for every observation, regardless of sensor readings or operating conditions. While this may seem trivial, it establishes a critical reference point. In highly imbalanced datasets like ours (96.7% non-failure cases), a naive model that always predicts the majority class achieves 96.7% accuracy. This demonstrates why accuracy is a misleading metric for imbalanced problems. The no-skill model catches zero failures, misses all 66 failures in the test set, and incurs the full 1,650,000 USD baseline cost. Any meaningful predictive model must outperform this baseline.
Model L1: Balanced Logistic Regression. The first predictive model applies logistic regression, a fundamental statistical classification technique. Logistic regression models the log-odds of failure as a linear combination of input features: air temperature, process temperature, rotational speed, torque, tool wear, and one-hot encoded machine type. To address class imbalance [10], we employ scikit-learn’s [4] class_weight=‘balanced’ parameter, which automatically adjusts the loss function to penalize misclassification of the minority class (failures) more heavily than the majority class (non-failures).
This approach uses the default decision threshold of 0.5 probability–if the model predicts a failure probability above 50%, it classifies the observation as a failure. Features are standardized to zero mean and unit variance to ensure numerical stability and proper weight scaling. The model is trained on 8,000 machines and evaluated on the held-out test set of 2,000 machines.
Model L2: Cost-Aware Logistic Regression. While Model L1 addresses class imbalance, it still treats all prediction errors equally. Model L2 enhances the logistic regression approach by optimizing the decision threshold explicitly for business costs rather than statistical balance [7].
After training the same logistic regression model as Model L1, we perform a grid search over possible probability thresholds from 0.01 to 0.99 to find the optimal threshold to minimize the total annual cost. For each candidate threshold, we generate predictions on the training set, calculate the resulting confusion matrix, apply our business cost model (5K USD for TP, 500 USD for FP, 25K USD for FN, 0 USD for TN), and compute total cost. The threshold that minimizes total training-set cost is selected as the final decision boundary. This cost-aware optimization explicitly encodes the 50-to-1 cost asymmetry between false negatives and false positives, shifting the decision boundary to favor higher recall even at the expense of increased false alarms.
Note that the threshold is learned on the training set and treated as a learned hyperparameter for test set evaluation.
Model L3: Cost-Aware Decision Tree. While logistic regression models linear relationships, many real-world failure mechanisms involve complex feature interactions. A heat dissipation failure, for example, may only occur when the temperature differential is low AND rotational speed is slow simultaneously - a conjunction that linear models struggle to capture.
Model L3 employs a decision tree classifier to learn non-linear decision boundaries through recursive feature splits. Decision trees naturally handle feature interactions by creating hierarchical rules: “If temperature differential < 8.6K, then if rotational speed < 1380 RPM, predict failure.” To prevent overfitting, we constrain the tree with hyperparameters: maximum depth of 5 levels, minimum 50 samples required to split a node, and minimum 20 samples required in each leaf node.
After training this regularized decision tree, we apply the same cost-aware threshold optimization as Model L2. For each probability threshold (decision trees can output class probabilities), we compute total business cost on the training set and select the threshold that minimizes cost. This combination of non-linear modeling and business-aligned optimization represents the state-of-the-art for correlation-based approaches.
3.2.3. Probabilistic and Causal Approaches
To explore whether fundamentally different modeling paradigms can improve upon correlation-based methods, we developed two advanced models incorporating uncertainty quantification and formal causal inference.
Model L4: Bayesian Logistic Regression. While standard logistic regression produces point estimates of failure probability, Bayesian approaches quantify uncertainty in those estimates. Model L4 implements Bayesian logistic regression using PyMC [5], a probabilistic programming framework.
The model architecture mirrors Model L1’s logistic regression structure but replaces point-estimated parameters with probability distributions. Each feature weight is assigned a Normal prior distribution with mean 0 and standard deviation 10, expressing weak prior belief that features have moderate effects. The intercept term receives a Normal(0, 10) prior. Rather than finding a single “best” set of parameters, we use Markov Chain Monte Carlo (MCMC) sampling to explore the posterior distribution of parameters given the data.
Specifically, we run the No-U-Turn Sampler (NUTS) with 500 posterior samples per chain, 300 tuning iterations, and 2 parallel chains, targeting an acceptance rate of 0.85. This produces 1,000 posterior samples representing plausible parameter values. For prediction, we average predicted probabilities across all posterior samples, yielding a probability estimate that incorporates parameter uncertainty.
After obtaining Bayesian predictions, we again optimize the decision threshold for business costs. The key advantage of the Bayesian approach is interpretability: we can quantify uncertainty in predictions and examine posterior distributions of feature weights to assess which sensors provide reliable predictive signal.
Model L5: Causal Model. All previous models (L0-L4) learn correlations in observational data. They identify that certain sensor patterns co-occur with failures but cannot distinguish causation from correlation. Model L5 takes a fundamentally different approach: constructing an explicit causal model and using formal causal inference methods to identify causal effects.
Figure 1. Causal directed acyclic graph (DAG) for CNC machine failures. The graph encodes domain knowledge about physical mechanisms: exogenous variables (sensor readings and machine type) affect endogenous variables (temperature differential, power, and overstrain) which directly cause machine failures.
The causal graph (Figure 1) encodes engineering principles governing CNC machine failure. The graph captures three physical mechanisms:
Thermal failure occurs when insufficient temperature differential (process minus ambient air) prevents heat dissipation, causing component degradation.
Power failure results when mechanical power (torque × rotational speed) exceeds electrical and mechanical capacity.
Overstrain failure reflects cumulative stress where tool wear × torque exceeds material limits.
The graph models these derived variables (temperature differential, power, overstrain) as intermediate nodes mediating sensor effects on failure. This structure reflects CNC physics: sensors combine through physical laws to create stress conditions exceeding thresholds, making causal relationships invariant to operational changes.
The methodology follows a five-step causal inference pipeline
  1. 1)
    Encode domain knowledge as a DAG (Figure 1) with nodes (sensor measurements, machine type, failure) and edges (causal relationships).
  2. 2)
    Define treatment variables: tool wear (replacement timing), torque (operational parameters), and rotational speed (machine settings). Temperature variables are outcomes, not treatments.
  3. 3)
    Apply Pearl’s do-calculus [2] to determine causal effect identifiability. The backdoor criterion [3] identifies adjustment sets blocking confounding paths.
  4. 4)
    Estimate causal effects by fitting linear regressions with identified adjustment sets, obtaining unbiased causal coefficients.
  5. 5)
    Create derived causal features: temperature differential (thermal stress), power (energy dissipation), and overstrain (cumulative mechanical stress) with clear physical interpretations.
Finally, we train a Gradient Boosting Classifier [12] on both the original sensor features and the derived causal features. Gradient boosting builds an ensemble of weak learners (decision trees) that collectively capture complex non-linear relationships. As with previous models, we optimize the decision threshold for business costs.
The critical advantage of this causal approach is that it provides both predictive performance and interpretability. When the model predicts a failure, we can trace the prediction back to specific causal pathways—overstrain exceeding safe limits, thermal stress from inadequate cooling, or power demand exceeding capacity. This transparency enables targeted interventions rather than generic maintenance.
Model L6: Rule-Based Causal Model. We also implement a rule set that mirrors the engineering heuristics described in the dataset documentation: failures are declared when any causal mechanism (tool wear band, insufficient heat dissipation, unsafe power level, or torque-driven overstrain) breaches its learned threshold. Random search tunes those thresholds to minimize cost. The approach provides perfect recall, but, as discussed later, it generates hundreds of precautionary inspections and is therefore best interpreted as an upper bound on achievable recall rather than a deployable policy.
Model L7: PyMC BART. Finally, we train a Bayesian Additive Regression Tree (BART) classifier using the PyMC-BART implementation [14][5]. BART is a non-parametric ensemble that automatically discovers non-linear combinations of features while delivering posterior uncertainty for every prediction. We feed BART the same causal feature set as L5 to test whether a flexible Bayesian learner can match causal performance without the explicit do-calculus pipeline. When we also let BART consume raw sensor readings and automatically generated interaction terms, it rediscovers the same latent constructs (temperature differential, power, overstrain), but the absence of structural constraints increases false positives by 232-3X. This sensitivity analysis confirms the intuition that automatic feature generation alone can approximate causal features, and yet graph-informed regularization is still required to keep operational costs low.
A summary of the modeling techniques is in the table Table 5.
Model Description Strategy
L0: No-Skill Baseline No-skill reference model Always predict “no failure”
L1: Balanced Logistic Regression Linear classifier with class balancing Standard logistic regression with balanced class weights
L2: Cost-Aware Logistic Regression Threshold-optimized logistic regression Same as L1 but with cost-optimized threshold
L3: Cost-Aware Decision Tree Non-linear decision model Hyperparameter-tuned decision tree with cost-aware threshold
L4: Bayesian Logistic Regression Probabilistic logistic model Bayesian inference with uncertainty quantification
L5: Causal Model Causal structural model Causal graph–based modeling
L6: Rule-Based Causal Model Hand-crafted causal rules Learned thresholds on mechanistic rules
L7: PyMC BART Model Bayesian tree ensemble Non-parametric sum-of-trees with causal features
Table 5. Model Summary Table
3.2.4. Validation Strategy
To ensure results are robust and not artifacts of a single train-test split, we implemented a comprehensive validation strategy. All models (L0 through L5) were trained and evaluated across five different random seeds (42, 43, 44, 45, and 46). For each seed, the dataset is randomly shuffled before the 80-20 stratified split, producing different training and test sets while maintaining the 3.3% failure rate in both partitions.
Performance metrics are computed independently for each seed, then averaged to produce final results. This multi-seed validation ensures that observed performance differences reflect genuine model capabilities rather than fortunate (or unfortunate) data partitioning. If a model performs well on one seed but poorly on others, this indicates instability and overfitting. Conversely, consistent performance across seeds demonstrates robust generalization.
Additionally, we examine the generalization gap; the difference between training-set performance and test-set performance. A large gap indicates overfitting: the model has memorized training data patterns that do not generalize. We report both training and test savings percentages, with the gap measured in percentage points.
4. Results and Analysis
Table 6 summarizes test-set performance for all six models, averaged across five random seeds. Each model was evaluated on 2,000 held-out machines containing 66 deterministic failures, with a baseline reactive maintenance cost of 1,650,000 USD.
Model Total Cost [USD] Savings [USD] Savings % Recall Precision F1 Score TP FP FN
L0: No-Skill 1,650,000 0 0.0% 0.0% 0.0% 0.000 0 0 66
L1: Balanced LogReg 743,000 907,000 55.0% 81.8% 13.5% 0.232 54 346 12
L2: Cost-Aware LogReg 794,000 856,000 51.9% 83.3% 10.1% 0.181 55 488 11
L3: Cost-Aware Tree 572,500 1,077,500 65.3% 87.9% 26.0% 0.401 58 165 8
L4: Bayesian LogReg 878,000 772,000 53.2% 81.8% 13.5% 0.232 54 346 12
L5: Causal 492,500 1,157,500 70.2% 87.9% 92.1% 0.899 58 5 8
L6: Rule-Based Causal 477,000 1,173,000 71.1% 100.0% 18.3% 0.310 66 294 0
L7: PyMC BART 541,000 1,109,000 67.2% 89.4% 29.4% 0.442 59 142 7
Table 6. Model performance, costs, and savings
These results reveal a clear performance hierarchy:
  • Model L0 provides zero value, establishing that any real predictive model must outperform the trivial solution of doing nothing.
  • Models L1, L2, and L4—all variations of logistic regression—cluster around a 51-55% cost reduction, saving between 772,000 USD and 907,000 USD annually.
  • Model L3, leveraging non-linear decision trees, achieves a 65.3% cost reduction with 1,077,500 USD in annual savings.
  • Model L5, using formal causal inference, achieves a 70.2% cost reduction with 1,157,500 USD in annual savings while keeping false alarms in the single digits, close to zero.
  • Model L6’s rule-based thresholds push recall to 100% and edge out L5 in total cost, but do so by triggering 294 inspections—nearly five times more than the cost-aware tree—which would overwhelm maintenance crews.
  • Model L7’s PyMC BART sits between L3 and L5: it matches their recall and yet generates more false positives because it lacks hard structural constraints, illustrating the limits of “feature generation only” strategies.
4.1. Detailed Model Analysis
L0: No-Skill Baseline. As expected, the no-skill model does not contribute anything. By always predicting “no failure,” it achieves the naive accuracy of 96.7% (equal to the base rate of non-failures) while catching zero actual failures. All 66 failures in the test set occur unexpectedly, each incurring the full 25,000 USD emergency cost, totaling 1,650,000 USD. This model serves only as the reference point against which all other approaches must be measured.
L1: Balanced Logistic Regression. The first predictive model demonstrates that even simple linear models can extract meaningful signal from sensor data. Model L1 catches 54 out of 66 failures (81.8% recall), preventing 1,350,000 USD in potential downtime costs. These 54 true positives cost 5,000 USD each in preventive maintenance, totaling 270,000 USD.
However, the model also generates 346 false positives—unnecessary maintenance interventions on healthy machines. At 500 USD each, these false alarms cost 173,000 USD in wasted labor. The 12 missed failures (false negatives) each incur the full 25,000 USD emergency cost, totaling 300,000 USD. Summing all components yields total cost of 743,000 USD, representing 907,000 USD in savings compared to reactive maintenance—a 55% cost reduction.
The model’s 13.5% precision indicates that only about 1 in 7 alarms corresponds to a real failure. While this may seem poor, it is acceptable given the cost structure: each real failure costs 25,000 USD while each false alarm costs only 500 USD. The model can afford 50 false alarms per true positive and still break even economically.
L2: Cost-Aware Logistic Regression. Model L2 applies the same logistic regression architecture as L1 but optimizes the decision threshold explicitly for business costs. The threshold optimization pushes the decision boundary lower, making the model more aggressive in predicting failures. This increases recall slightly to 83.3% (55 out of 66 failures caught) but generates substantially more false alarms—488 instead of 346.
The cost breakdown reveals why this strategy backfires. The 55 true positives cost 275,000 USD, the 488 false positives cost 244,000 USD, and the 11 false negatives cost 275,000 USD, totaling 794,000 USD. While the model catches one additional failure compared to L1, it generates 142 additional false alarms, adding 71,000 USD in unnecessary intervention costs. Net savings of 856,000 USD are actually 51,000 USD lower than L1.
This counterintuitive result illustrates a subtle point: threshold optimization on the training set can lead to overfitting of the decision boundary. The optimal threshold for training data may be too aggressive for test data if the model’s probability calibration is poor [11]. Model L2’s probabilities are not well-calibrated—many predicted “high risk” machines do not actually fail—so lowering the threshold excessively inflates false alarms.
L3: Cost-Aware Decision Tree. Model L3 represents the best correlation-based approach by combining non-linear modeling with constrained complexity. The decision tree architecture naturally captures feature interactions such as “heat dissipation failure occurs when temperature differential is low AND rotational speed is low.” Regularization hyperparameters (maximum depth 5, minimum 50 samples per split, minimum 20 samples per leaf) prevent the tree from memorizing training data noise.
This model catches 58 out of 66 failures (87.9% recall), a substantial improvement over L1’s 54. Critically, it does so with only 165 false alarms—less than half of L1’s 346 false positives. This superior precision (26.0%, meaning about 1 in 4 alarms is real) reflects the model’s ability to identify genuine failure patterns rather than overfitting to spurious correlations.
The cost breakdown is favorable: 58 true positives at 5,000 USD each cost 290,000 USD, 165 false positives at 500 USD each cost 82,500 USD, and 8 missed failures at 25,000 USD each cost 200,000 USD. Total cost of 572,500 USD yields 1,077,500 USD in annual savings—a 65.3% cost reduction. This represents the ceiling of what correlation-based machine learning can achieve on this problem.
L4: Bayesian Logistic Regression. Model L4 applies Bayesian inference to quantify prediction uncertainty, but this additional sophistication does not translate to improved test-set performance. The model achieves 81.8% recall and 13.5% precision, identical to L1’s results. Total cost of 878,000 USD yields 772,000 USD in savings—better than reactive maintenance but worse than simple balanced logistic regression (L1).
The modest performance suggests that parameter uncertainty, while theoretically valuable, does not substantially improve predictions when the underlying model class (linear logistic regression) has limited expressiveness. The Bayesian framework provides interpretability—we can examine posterior distributions to assess which features have reliably non-zero effects—but does not overcome the fundamental limitation that linear models cannot capture the complex non-linear interactions present in failure mechanisms.
The computational cost of Bayesian inference (several minutes for MCMC sampling versus seconds for standard logistic regression) further reduces practical appeal unless the uncertainty quantification itself is valuable for decision-making. For this predictive maintenance application, actionable predictions matter more than prediction intervals.
L5: Causal Model. Model L5 achieves the highest performance across all metrics, demonstrating the value of causal reasoning. The model catches 58 out of 66 failures (87.9% recall), matching Model L3’s recall. However, it does so with only 5 false positives compared to L3’s 165—a dramatic 97% reduction in false alarms. Precision of 92.1% means that more than 9 out of 10 alarms correspond to real failures.
This exceptional precision-recall balance translates to substantial cost savings. The 58 true positives cost 290,000 USD, the mere 5 false positives cost only 2,500 USD (negligible), and the 8 missed failures cost 200,000 USD. Total cost of 492,500 USD yields 1,157,500 USD in annual savings—a 70.2% cost reduction and the highest financial value among all models.
The model’s superiority stems from its causal architecture. By encoding domain knowledge in the causal graph and deriving physically meaningful features (temperature differential for thermal stress, power for energy dissipation, overstrain for mechanical stress), the model learns relationships that generalize robustly. When the model predicts a failure, it does so because a specific causal mechanism has been activated—not because of a spurious correlation in the training data.
For example, a prediction might be traced to high overstrain: tool wear has reached 215 minutes while torque is 56 Nm, producing an overstrain index of 12,040 that exceeds the safe threshold of 11,000. This level of interpretability enables targeted interventions: the maintenance team can replace the tool before reaching the dangerous threshold, addressing the root cause directly.
L6: Rule-Based Causal Model. The threshold-based rule system provides a useful upper bound on recall. By declaring a failure whenever any mechanism is even close to its safety margin, the rules catch all 66 deterministic failures (100% recall) and reduce total cost to 477,000 USD. However, the policy produces 294 false alarms—orders of magnitude more than the five false alarms produced by L5. In practice that would translate into hundreds of unnecessary stoppages per week. We therefore treat L6 as a “safety-first” bound that validates the physics of the rules, not as a candidate deployment policy.
L7: PyMC BART Model. BART delivers 1,109,000 USD in savings (67.2%) with recall of 89.4%, but its precision drops to 29.4% (142 false positives). This outcome answers potential questions around automatic feature generation: even a powerful Bayesian ensemble that can discover complex interactions still benefits from graph-regularized features. Without explicit causal constraints the model develops conservative decision boundaries, trading cost savings for a larger inspection queue.
4.1.1. Comparative Analysis: Causal vs. Correlation
The performance gap between Model L5 (causal) and Model L3 (best correlation) quantifies the value of causal reasoning. Model L5 saves an additional 80,000 USD annually (7.4% improvement in savings percentage) while generating 160 fewer false alarms—a 97% reduction in operational disruption. L6 shows that we can squeeze a few more dollars out of the problem by firing every possible alarm, but at the cost of 294 inspections; L5 reaches nearly the same savings with almost no false positives.
This improvement cannot be attributed to model complexity. L3 is a single, carefully regularized decision tree rather than a gradient-boosted ensemble, whereas L5 couples do-calculus adjustments with gradient boosting and L7 adds a Bayesian additive regression tree. When we train powerful ensembles (gradient boosting and BART) on raw features without causal guidance, they rediscover similar splits yet inflate the false-positive rate. The difference lies in the features themselves. Model L3 learns from raw sensor readings, discovering correlations like “high torque readings are associated with failure.” Model L5 learns from causally derived features, understanding mechanisms like “high torque combined with worn tools causes mechanical overstrain failure,” while Model L7 demonstrates that even a flexible Bayesian ensemble still benefits from causal priors.
When operating conditions shift slightly—perhaps a new batch of machines has different temperature baselines, or a process change alters typical torque ranges—Model L3’s correlations may degrade, while Model L5’s causal relationships remain valid. The physics of mechanical overstrain do not change when ambient temperature drifts, whereas the correlation between specific torque values and failure does.
This distinction has profound implications for model deployment and longevity. Model L3 requires periodic retraining as data distributions shift. Model L5’s causal structure provides inherent robustness to distribution changes, as causal relationships are invariant properties of the system rather than contingent patterns in historical data.
4.1.2. Generalization and Overfitting Assessment
To assess whether observed performance reflects genuine model capability or lucky train-test splits, we examine the generalization gap—the difference between training performance and test performance. Table 7 shows training savings percentage, test savings percentage, and the gap in percentage points for predictive models.
Model Train Savings % Test Savings % Gap (%)
L1: Balanced LogReg 56.9% 55.0% 1.97
L2: Cost-Aware LogReg 55.8% 51.9% 3.93
L3: Cost-Aware Tree 69.4% 65.3% 4.11
L4: Bayesian LogReg 56.9% 53.2% 3.70
L5: Causal 72.8% 70.2% 2.60
L6: Rule-Based Causal 73.5% 71.1% 2.40
L7: PyMC BART 69.1% 67.2% 1.87
Table 7. Train–test savings comparison. Gap represents the difference between training and test savings percentages; smaller gaps indicate better generalization.
All models exhibit modest generalization gaps below 5 percentage points, which we set as an acceptable bound because plant operators typically budget that amount of degradation when deploying predictive maintenance models across production lines [8]. The multi-seed experiment (five independent splits) confirmed that variance stays within this tolerance without any additional regularization. Model L3 shows the largest gap at 4.11 percentage points, suggesting slight over-optimization to training data, yet its test performance remains the highest among correlation-based approaches.
Logistic regression (L1) records the smallest gap at 1.97 percentage points, with BART close behind at 1.87 percentage points. Model L5’s 2.60 point gap is only marginally larger while delivering far superior cost metrics, reinforcing the hypothesis that causal features generalize more robustly than raw correlations even if they do not literally yield the lowest generalization gap.
These findings are reinforced by multi-seed validation. Across five different random splits, model rankings remained completely stable: L5 outperformed L3, which outperformed L1, which outperformed L2, which outperformed L0. This consistency demonstrates that observed performance differences reflect true model capabilities rather than chance variations in data partitioning.
5. Conclusion
This study evaluated six predictive maintenance models on a CNC machine failure dataset, comparing traditional correlation-based machine learning approaches with formal causal inference methods. Results indicate that the causal AI approach (Model L5) achieved superior performance on business-relevant metrics, delivering 1,157,500 USD in annual cost savings (70.2% reduction) while maintaining high recall (87.9%) and exceptional precision (92.1%). The causal model outperformed the best correlation-based decision tree (Model L3) by 80,000 USD annually while reducing false alarms by 97%.
This work demonstrates several important findings for predictive maintenance applications:
  1. (1)
    Business-aligned optimization: Models optimized explicitly for business costs (accounting for the 50:1 cost asymmetry between false negatives and false positives) substantially outperform accuracy-optimized approaches.
  2. (2)
    Causal superiority: When domain knowledge enables construction of causal graphs, formal causal inference methods can provide both superior predictive performance and interpretable root-cause explanations for predictions.
  3. (3)
    Bayesian ensembles still benefit from structure: The PyMC BART model validates that automatic feature generation can approach causal performance, but without graph-informed constraints, it issues substantially more false positives.
  4. (4)
    Robust generalization: Causally-derived features demonstrate stronger generalization (2.6 percentage point train-test gap) compared to raw sensor correlations (4.1 percentage points for the best correlation model), consistent with theoretical expectations that causal relationships are invariant under distribution shifts.
  5. (5)
    Operational benefits: The dramatic reduction in false alarms (97% fewer than correlation-based approaches) minimizes operational disruption while maintaining safety through high recall.
Several limitations of this study should be acknowledged:
  1. (1)
    Dataset scope: The analysis uses a single dataset of 10,000 CNC machines. Generalization to other machine types, manufacturing environments, or operational conditions requires additional validation.
  2. (2)
    Causal graph construction: The causal DAG was manually constructed based on domain knowledge of CNC machine failure mechanisms. In domains where such expertise is unavailable, causal discovery algorithms could be explored, though they introduce additional challenges.
  3. (3)
    Cost estimates: The business costs (25,000 USD for failures, 5,000 USD for preventive maintenance, 500 USD for false alarms) are representative estimates. Actual costs vary by industry, machine criticality, and operational context.
  4. (4)
    Static analysis: This study evaluates models on historical data without real-world deployment. Practical implementation would require monitoring for distribution drift, model degradation, and integration with existing maintenance management systems.
  5. (5)
    Class imbalance: The 3.3% failure rate, while realistic for well-maintained equipment, limits the number of positive examples available for model training. Larger datasets with more failure observations could improve model performance.
Future research directions include: (1) validation across multiple manufacturing domains and equipment types; (2) investigation of causal discovery algorithms to automate DAG construction; (3) development of online learning approaches for continual model adaptation; (4) integration with IoT sensor networks and SCADA systems for real-time deployment; and (5) extension to multi-component systems with cascading failure modes.