### v2

# A Benchmark of Causal vs. Correlation AI for Predictive Maintenance

Abstract.
Predictive maintenance in manufacturing environments presents a challenging optimization problem characterized by extreme cost asymmetry, where missed failures incur costs roughly fifty times higher than false alarms. Conventional machine learning approaches typically optimize statistical accuracy metrics that do not reflect this operational reality and cannot reliably distinguish causal relationships from spurious correlations. This study benchmarks eight predictive models, ranging from baseline statistical approaches to Bayesian structural causal methods, on a dataset of 10,000 CNC machines with a 3.3 percent failure prevalence. While ensemble correlation-based models such as Random Forest (L4) achieve the highest raw cost savings (70.8 percent reduction), the Bayesian Structural Causal Model (L7) delivers competitive financial performance (66.4 percent cost reduction) with an inherent ability of failure attribution, which correlation-based models do not readily provide. The model achieves perfect attribution for HDF, PWF, and OSF failure types. These results suggest that causal methods, when combined with domain knowledge and Bayesian inference, offer a potentially favorable trade-off between predictive performance and operational interpretability in predictive maintenance applications.
[…]
1. Contributions
This study makes the following contributions to predictive maintenance optimization:
  • Systematic benchmark of causal vs. correlation-based models: We evaluate eight models ranging from basic baselines to advanced Bayesian structural causal methods on a dataset of 10,000 CNC machines from the UCI Machine Learning Repository [4]. This study provides a comprehensive comparison of formal causal inference against traditional ML approaches for predictive maintenance.
  • Business-aligned optimization framework: We demonstrate that models optimized explicitly for business costs (accounting for the 50:1 cost asymmetry between false negatives and false positives) substantially outperform accuracy-optimized approaches, which highlights the critical importance of cost-sensitive learning in industrial applications.
  • Failure attribution without attribution labels: The Bayesian SCM (L7) attributes each predicted failure to its most likely physical mechanism (TWF, HDF, PWF, or OSF) without requiring failure-type labels during training. This allows for maintenance interventions with a precise focus on the identified root cause.
  • Competitive performance with interpretability: The Bayesian SCM achieves a 66.4% cost reduction, a competitive result compared with the state-of-the-art correlation-based ensemble models. Together with the failure attribution capability, this shows that causal AI can deliver a favorable trade-off between predictive performance and interpretability.
  • Ensemble methods as a performance ceiling: Tree-based ensemble models (Random Forest, AdaBoost, LightGBM) consistently outperform single models including the decision tree, which mirrors the true data-generating process for the synthetic CNC data set. This proves the effectiveness of ensemble methods in reducing variance and preventing overfitting.
2. Introduction
Manufacturing faces a maintenance dilemma. Reactive maintenance leads to costly downtime, while overly cautious preventive maintenance wastes resources on healthy equipment. The core challenge is an extreme cost asymmetry across industrial equipment: a missed failure can cost tens of thousands of dollars (unplanned downtime, emergency repairs, production losses), while a false alarm may cost only hundreds of dollars (scheduled inspection, standard parts). For the CNC machines analyzed in this study, we use representative costs of 25,000 USD for missed failures and 500 USD for false alarms, reflecting typical manufacturing industry values where the cost ratio between unplanned downtime and preventive maintenance ranges from 20:1 to 100:1 depending on equipment criticality. Table 1 summarizes the four possible prediction outcomes, their associated costs, and business impact.
Outcome Description Cost Business Impact
True Positive (TP) Correctly predict failure 5,000 USD Scheduled repair prevents downtime
False Positive (FP) Incorrectly predict failure 500 USD Wasted inspection/parts
False Negative (FN) Miss actual failure 25,000 USD Emergency downtime + repairs
True Negative (TN) Correctly predict no failure 0 USD Normal production
Table 1. Prediction outcomes, costs, and business impact. Cost estimates reflect typical manufacturing industry values where unplanned downtime incurs significant production losses, emergency labor, and expedited parts costs, while preventive maintenance requires only scheduled labor and standard parts procurement [8].
Standard machine learning models optimize for statistical accuracy metrics such as accuracy, precision, recall, F1-score, and AUC-ROC, which either ignore class imbalance or treat all prediction errors equally. Accuracy is misleading in imbalanced datasets. In fact, a model that always predicts “no failure” achieves 96.7% accuracy on our data while providing no business value. Precision focuses on minimizing false alarms but may miss critical failures, while recall prioritizes catching all failures but may trigger hundreds of unnecessary inspections. F1-score balances precision and recall through their harmonic mean, and AUC-ROC evaluates the model’s ranking ability across all possible thresholds, but neither metric accounts for the actual business costs of different error types. All these metrics fail to capture the 50:1 cost ratio between false negatives and false positives that is typical for manufacturing equipment, where unplanned downtime costs far exceed scheduled maintenance [8]. Furthermore, correlation-based models identify patterns that co-occur with failures but cannot distinguish causation from spurious relationships. This limits their utility to a binary alarm with no actionable diagnosis.
Our goal is to determine if Causal AI can deliver competitive business value relative to state-of-the-art correlation-based ML while also offering reliable insights into the failure root causes in predictive maintenance scenarios.
3. Dataset and Experimental Design
Our analysis utilizes the UCI Machine Learning Repository’s CNC Machine Predictive Maintenance dataset, which is a synthetic dataset designed in [4] to reflect real-world predictive maintenance scenarios. The dataset contains 10,000 machines and for each machine six sensor measurements as described in Table 2:
Feature Description
Air temperature Measured in Kelvin
Process temperature Measured in Kelvin
Rotational speed Revolutions per minute
Torque Newton-meters
Cumulative tool wear Minutes of operation
Machine type classification Low, Medium, or High specification
Table 2. Feature descriptions
A snippet of the data is shown in Table 3.
UDI Product ID Type Air temp [K] Process temp [K] Rot. speed [rpm] Torque [Nm] Tool wear [min] Machine failure TWF HDF PWF OSF RNF
1 M14860 M 298.1 308.6 1551 42.8 0 0 0 0 0 0 0
2 L47181 L 298.2 308.7 1408 46.3 3 0 0 0 0 0 0
3 L47182 L 298.1 308.5 1498 49.4 5 0 0 0 0 0 0
4 L47183 L 298.2 308.6 1433 39.5 7 0 0 0 0 0 0
5 L47184 L 298.2 308.7 1408 40.0 9 0 0 0 0 0 0
Table 3. Dataset snippet showing the first five rows of the CNC machine predictive maintenance dataset.
The dataset documents five distinct failure modes that can occur in CNC machines, as summarized in Table 4:
Symbol Failure Type Description Instances
TWF Tool Wear Failure Tool replacement or failure at randomly selected wear time 46
HDF Heat Dissipation Failure Process failure due to temperature differential (process minus air) 115
PWF Power Failure Process failure due to mechanical power (torque $\times$ rotational speed in rad/s) 95
OSF Overstrain Failure Process failure due to tool wear $\times$ torque exceeds threshold 98
RNF Random Failure 0.1% probability of failure regardless of process parameters (unpredictable) 19
Table 4. Failure types with symbols, descriptions, and instance counts. Definitions follow the dataset documentation [4].
3.1. Target Definition and Rationale
For this analysis, we focus exclusively on predictable failures—those failure modes that can theoretically be predicted from sensor measurements and operating conditions. Specifically, our prediction target combines Tool Wear Failure, Heat Dissipation Failure, Power Failure, and Overstrain Failure, while explicitly excluding Random Failure events. Random failures are defined in the dataset as having a fixed 0.1% probability of occurrence regardless of process parameters. They are statistically independent of all sensor readings and represent truly unpredictable noise (only 19 instances in 10,000 data points). It is worth noting that while TWF is not excluded, it is largely nondeterministic. In fact, the data generation process of TWF contains an irreducible stochastic component. Readers may refer to [5] or Section 4.1.1 for more details.
After applying this filter, our dataset contains 9,981 total observations. We employ a stratified 80-20 train-test split that preserves the class distribution in both sets. The test set comprises 2,000 machines among which 66 experience deterministic failures (3.3% failure rate). This severe class imbalance poses significant challenges for standard machine learning approaches and underscores why accuracy-based metrics are inappropriate for evaluation.
  • Size: 10,000 machines with sensor readings
  • Features: Air temperature, process temperature, rotational speed, torque, tool wear, machine type (L/M/H)
  • Target: Deterministic failures excluding random failures (TWF + HDF + PWF + OSF)
  • Test Set: 2,000 machines (20% stratified split) with 66 failures (3.3% prevalence)
  • Baseline Cost: 1,650,000 USD (reactive maintenance: 66 failures $\times$ 25,000 USD)
3.2. Methodology
3.2.1. Success Metrics
The metrics we consider are the following:
  1. 1.
    Total annual cost: The business objective is to minimize the total cost of machine maintenance
  2. 2.
    Savings: Dollar reduction vs. reactive maintenance baseline
  3. 3.
    Recall: Percentage of actual failures caught
  4. 4.
    Precision: Percentage of alarms that are real failures
  5. 5.
    Generalization: Train vs. test performance gap
3.2.2. Correlation-Based Models (L0–L3)
We start by investigating four correlation-based models, ranging from a no-skill baseline through logistic regression variants to a decision tree.
Model L0: No-Skill Baseline. The no-skill model serves as the minimal performance floor. This model predicts “no failure” for every observation, regardless of sensor readings or operating conditions. While this may seem trivial, it establishes a critical reference point. In highly imbalanced datasets like ours (96.7% non-failure cases), a naive model that always predicts the majority class achieves high accuracy. This demonstrates why accuracy is a misleading metric for imbalanced problems. The no-skill model catches zero failures, misses all 66 failures in the test set, and incurs the full 1,650,000 USD baseline cost. Any meaningful predictive model must outperform this baseline.
Model L1: Balanced Logistic Regression. The first predictive model applies logistic regression, which is a fundamental classification technique. Logistic regression models the log-odds of failure as a linear combination of input features: air temperature, process temperature, rotational speed, torque, tool wear, and one-hot encoded machine type. To address class imbalance [2], we employ scikit-learn’s [6] class_weight=‘balanced’ parameter, which automatically adjusts the loss function to penalize misclassification of the minority class (failures) more heavily than the majority class (non-failures).
This approach uses the default decision threshold of 0.5 probability–if the model predicts a failure probability above 50%, it classifies the observation as a failure. Features are standardized to zero mean and unit variance to ensure numerical stability and proper weight scaling. The model is trained on 8,000 machines and evaluated on the held-out test set of 2,000 machines.
Model L2: Cost-Aware Logistic Regression. While Model L1 addresses class imbalance, it still treats all prediction errors equally. Model L2 enhances the logistic regression approach by optimizing the decision threshold explicitly for business costs rather than statistical balance [1].
After training the same logistic regression model as Model L1, we perform a grid search over possible probability thresholds from 0.01 to 0.99 to find the optimal threshold to minimize the total annual cost. For each candidate threshold, we generate predictions on the training set, calculate the resulting confusion matrix, apply our business cost model (5K USD for TP, 500 USD for FP, 25K USD for FN, 0 USD for TN), and compute total cost. The threshold that minimizes total training-set cost is selected as the final decision boundary. This cost-aware optimization explicitly encodes the 50-to-1 cost asymmetry between false negatives and false positives. It encourages the model to shift the decision boundary to favor higher recall even at the expense of increased false alarms. Note that the threshold is learned on the training set and treated as a learned hyperparameter for test set evaluation.
Model L3: Cost-Aware Decision Tree. While logistic regression models linear relationships, many real-world failure mechanisms involve complex feature interactions. A heat dissipation failure, for example, may only occur when the temperature differential is low AND rotational speed is slow simultaneously. This presents a conjunction that linear models struggle to capture.
Model L3 employs a decision tree classifier to learn non-linear decision boundaries through recursive feature splits. Decision trees naturally handle feature interactions by creating hierarchical rules: “If temperature differential < 8.6K, then if rotational speed < 1380 RPM, predict failure.” To prevent overfitting, we constrain the tree with hyperparameters: maximum depth of 5 levels, minimum 50 samples required to split a node, and minimum 20 samples required in each leaf node.
After training this regularized decision tree, we apply the same cost-aware threshold optimization as Model L2. For each probability threshold (decision trees can output class probabilities), we compute total business cost on the training set and select the threshold that minimizes cost. This combination of non-linear modeling and business-aligned optimization represents the state-of-the-art for correlation-based approaches.
3.2.3. Ensemble Models (L4–L6)
Model L4: Random Forest. While the decision tree model is akin to the true data generating process of this dataset, in practice a single tree often suffers from high variance. A small change in training data can yield drastically different tree structures. Thus we consider the Random Forest model, which addresses this limitation via ensembling. A Random Forest model builds multiple decision trees and aggregates their results by averaging their predictions. This aggregation method reduces variance compared to using a single model, and helps prevent overfitting. In fact, during training we employed the following configuration:
  1. (1)
    Bootstrap sampling: Each tree is trained on a bootstrap sample, meaning that we draw $N$ training samples with replacement from the original $N$ samples. Each tree sees approximately 63% unique samples.
  2. (2)
    Random Feature Selection: We randomly select $\sqrt{7}\approx 3$ features as candidates when splitting a node. This prevents strong features from dominating every tree’s upper splits and encourages discovering diverse patterns.
  3. (3)
    Regularization Constraints: We require at least 10 samples to split a node. This helps prevent overfitting.
The Random Forest model is implemented using the scikit-learn library. After obtaining ensemble probability predictions, we apply cost-aware threshold optimization as in Models L2--L3 to minimize total business cost.
Model L5: AdaBoost. Now we turn to boosting algorithms. We first consider AdaBoost, which is a foundational boosting algorithm that demonstrates the strength of combined weak learners. AdaBoost trains the weak learners on a weighted version of the training data, where the weight on incorrectly classified samples increases exponentially after each iteration.
We implement AdaBoost using scikit-learn’s AdaBoostClassifier with decision stumps (depth-1 trees) as base learners. We train 50 decision stumps with learning rate 1.0. Each stump makes a binary decision based on one feature. The ensemble prediction is a weighted average over all stumps.
Finally, we apply the cost-aware threshold optimization as in Models L2--L3.
Model L6: Light Gradient Boosting Machine. As the final model in the ensemble class, we consider LightGBM, which represents the modern evolution of gradient boosting algorithms. It greatly reduces training time compared to earlier gradient boosting algorithms (e.g. XGBoost, CatBoost) while maintaining or improving accuracy. For more details regarding the innovations of LightGBM, the reader may refer to [3].
We implement LightGBM using the lightgbm Python library with 100 boosting iterations, maximum tree depth of 3, learning rate of 0.1, and 10 maximum leaves per tree. Cost-aware threshold optimization is applied at the end of the pipeline.
3.2.4. Causal Model
Finally, we consider a causal model that utilizes domain-specific failure mechanisms. This structure allows for both prediction and failure attribution, i.e. telling which failure mode (TWF, HDF, PWF, or OSF) is responsible.
Model L7: Bayesian Structural Causal Model (SCM). Unlike correlation models that learn $P(Y|X)$ by fitting weights to all available features, this model encodes the specific physical mechanisms of failure into a structural graph. We explicitly model four independent failure mechanisms that combine via a logical “Noisy-OR” gate to produce the final machine failure.
We define the system using a set of structural equations implemented in PyMC [7]. We standardized all input features $X$. In the following, we let $\sigma(z)=1/(1+e^{-z})$ be the sigmoid function.
1. Tool Wear Failure (TWF) Mechanism: This mechanism depends exclusively on the tool wear duration. We model the probability $p_{twf}$ as a logistic function of tool wear ($x_{tw}$), excluding all other variables to enforce causal independence:
(1) $p_{twf}=\sigma(\alpha_{twf}+\beta_{twf}\cdot x_{tw}).$
2. Heat Dissipation Failure (HDF) Mechanism: Physical heat failure occurs when the heat cannot be dissipated, which happens when the temperature difference ($x_{\Delta T}$) is low AND rotational speed ($\omega$) is low. We model this using an explicit interaction term:
(2) $p_{hdf}=\sigma(\alpha_{hdf}+\beta_{temp}x_{\Delta T}+\beta_{rot}\omega+\beta_{int}(x_{\Delta T}\cdot\omega)).$
3. Power Failure (PWF) Mechanism: Power failure occurs if the power ($P=\tau\cdot\omega$) is too low or too high. To capture this non-monotonic U-shaped relationship, we introduce a quadratic term $P^{2}$:
(3) $p_{pwf}=\sigma(\alpha_{pwf}+\beta_{p}P+\beta_{p^{2}}P^{2}).$
4. Overstrain Failure (OSF) Mechanism: Overstrain depends on the product of Tool Wear and Torque ($S=x_{tw}\cdot\tau$). The failure threshold depends on the machine quality type ($L,M,H$). We model this using dummy variables for Low ($\mathbb{I}_{L}$) and Medium ($\mathbb{I}_{M}$) types:
(4) $p_{osf}=\sigma(\alpha_{osf}+\beta_{osf}S+\beta_{L}\mathbb{I}_{L}+\beta_{M}\mathbb{I}_{M}).$
5. System Failure (Noisy-OR): Finally, we combine these independent probabilities using a Noisy-OR gate. This logic implies that the machine survives only if it survives all individual failure mechanisms:
(5) $P(\text{Failure})=1-\left[(1-p_{twf})(1-p_{hdf})(1-p_{pwf})(1-p_{osf})\right].$
We place weakly informative Normal priors on all $\alpha$ and $\beta$ parameters (e.g., $\mathcal{N}(0,1)$) and use NUTS sampling to estimate the posterior distributions.
A summary of the modeling techniques is in Table 5.
Model Description Strategy
L0: No-Skill Baseline No-skill reference model Always predict “no failure”
L1: Balanced Logistic Regression Linear classifier with class balancing Standard logistic regression with balanced class weights
L2: Cost-Aware Logistic Regression Threshold-optimized logistic regression Same as L1 but with cost-optimized threshold
L3: Cost-Aware Decision Tree Non-linear decision model Hyperparameter-tuned decision tree with cost-aware threshold
L4: Random Forest Decision tree ensemble Average the predictions of multiple trees
L5: AdaBoost Boosted weak learner ensemble Sequentially reweight misclassified samples
L6: LightGBM Gradient boosting machine Histogram-based leaf-wise tree growth with boosting
L7: Bayesian SCM Bayesian structural causal model Explicit failure mechanisms with Noisy-OR combination
Table 5. Model Summary Table
3.2.5. Validation Strategy
To ensure results are robust and not artifacts of a single train-test split, we implemented a comprehensive validation strategy. All models (L0 through L7) were trained and evaluated across five different random seeds (42, 43, 44, 45, and 46). For each seed, the dataset is randomly shuffled before the 80-20 stratified split, producing different training and test sets while maintaining the 3.3% failure rate in both partitions.
Performance metrics are computed independently for each seed, then averaged to produce final results. This multi-seed validation ensures that observed performance differences reflect genuine model capabilities rather than fortunate (or unfortunate) data partitioning. If a model performs well on one seed but poorly on others, this indicates instability and overfitting. Conversely, consistent performance across seeds demonstrates robust generalization.
Additionally, we examine the generalization gap; the difference between training-set performance and test-set performance. A large gap indicates overfitting: the model has memorized training data patterns that do not generalize. We report both training and test savings percentages, with the gap measured in percentage points. A test performance within 5 percentage points of training performance we deem acceptable (following [8]).
4. Results and Analysis
Table 6 summarizes test-set performance for all models, averaged across five random seeds. Each model was evaluated on 2,000 held-out machines containing 66 deterministic failures, with a baseline reactive maintenance cost of 1,650,000 USD.
Model Total Cost [USD] Savings [USD] Savings % Recall Precision F1 Score TP FP FN
L0: No-Skill 1,650,000 0 0.0% 0.0% 0.0% 0.000 0 0 66
L1: Balanced LogReg 796,600 853,400 51.7% 77.3% 13.3% 0.227 51 333 15
L2: Cost-Aware LogReg 813,000 837,000 50.7% 82.4% 9.8% 0.175 54 502 11
L3: Cost-Aware Tree 576,900 1,073,100 65.0% 90.0% 23.6% 0.364 59 229 6
L4: Random Forest 481,500 1,168,500 70.8% 96.4% 23.6% 0.379 64 207 2
L5: AdaBoost 531,500 1,118,500 67.8% 95.2% 18.9% 0.314 63 275 3
L6: LightGBM 510,900 1,139,100 69.0% 91.5% 31.2% 0.364 60 138 6
L7: Bayesian SCM 554,900 1,095,100 66.4% 90.3% 26.2% 0.395 59 193 6
Table 6. Model performance, costs, and savings, averaged over validation splits. TP, FP, and FN counts are rounded to the nearest integer.
These results reveal a clear performance hierarchy:
  • Model L0 provides zero value, establishing that any real predictive model must outperform the trivial solution of doing nothing.
  • Models L1 and L2—variations of logistic regression—cluster around a 50–52% cost reduction, saving between 837,000 USD and 853,400 USD annually.
  • Model L3, leveraging non-linear decision trees, achieves a 65.0% cost reduction with 1,073,100 USD in annual savings.
  • Model L4 (Random Forest) achieves the best performance among all models with a 70.8% cost reduction and 1,168,500 USD in annual savings and the highest recall (96.4%), catching 64 out of 66 failures with only 2 missed.
  • Model L5 (AdaBoost) achieves a 67.8% cost reduction with 1,118,500 USD in annual savings and 95.2% recall (63 of 66 failures caught).
  • Model L6 (LightGBM) achieves a 69.0% cost reduction with 1,139,100 USD in annual savings. Though the model predicted considerably fewer false positives compared to random forest, its lower recall still results in slightly higher costs due to missed failures.
  • Model L7 (Bayesian SCM) achieves a 66.4% cost reduction with 1,095,100 USD in annual savings and 90.3% recall. While this places it behind the best ensemble models on raw cost metrics, it is the only model capable of attributing predicted failures to specific types, as discussed in Section 4.1.1.
4.1. Detailed Model Analysis
The no-skill baseline (L0) catches zero failures and incurs the full 1,650,000 USD reactive maintenance cost. This serves as baseline for evaluating other models. Logistic regression models and their variants (L1, L2) resulted in 47–55% cost reduction relative to the baseline. Their suboptimal results are not surprising considering the nonlinear true data generating process and the linear decision boundary these models use to make predictions.
The non-linear tree-based correlation-based models (L3, L4, L5, L6) perform substantially better with 65–71% cost reduction. Random Forest (L4) had the best performance as it misses only 2 failures with a reasonable number of false positives. It is worth noting that the underlying data generating process of this dataset is itself a decision tree. Yet ensemble models, even equipped with a weak learner like AdaBoost, outperform a single decision tree. This demonstrates the strength of ensemble methods as they aggregate multiple estimators and can overcome the single tree’s tendency to overfit. On the other hand, a single decision tree has better interpretability since one can trace the decision path down the tree to infer the failure mode. This property is lost in ensemble models, in exchange for superior accuracy and variance reduction.
The causal model, Bayesian SCM (L7), achieves a 66.4% cost reduction and the highest recall among non-ensemble models (90.3%). Compared to the best of the correlation-based models, the causal model suffers a modest performance loss in exchange for stronger interpretability. That is, when the model predicts a failure, it simultaneously attributes that failure to the most likely causes. This allows the maintenance team to address the specific failure mechanism identified by the model, which can be a favorable trade-off in real-world maintenance settings where understanding why a machine will fail is as important as knowing that it will.
4.1.1. Failure Attribution
Among the correlation-based models, the decision tree offers the strongest level of interpretability. One may trace the decision path, record the features that drive the classification and infer the failure types. This, however, requires additional effort to perform the backtracing. Other correlation-based models like ensemble methods present a deeper challenge, as the reasoning is masked behind the aggregation of many weak learners. The causal model, with the domain knowledge encoded in the structure equations, has the inherent ability to attribute failure without needing interpretation efforts.
We examine the quality of failure attribution reported by the Bayesian SCM. Recall that the model was trained using only binary machine failure labels without knowing specific failure types (TWF, HDF, PWF, or OSF). Inside the model it uses structural equations to estimate the probability of each failure mode. When the model predicts a failure, we attribute the failure to the type with the greatest probability. We evaluate the attribution results against the ground truth failure type labels available in the dataset. Table 7 reports the number of ground truth failures in the test set (seed 42), the number detected by the model, and the number correctly attributed among those detected.
Failure Type Ground Truth Detected Correctly Attributed
TWF 7 3 1
HDF 25 25 25
PWF 15 15 15
OSF 22 22 22
Table 7. Per-type failure attribution performance on the test set (seed 42). Attribution analysis is reported for a single representative seed as failure type labels were not recorded in the averaged results.
For those detected failures, the Bayesian SCM achieves 100% correct attribution for HDF, PWF, and OSF types. This encouraging result shows that the structural equations successfully captured the physical mechanisms of each failure type, especially considering that the model was never given the failure type labels during training. TWF attribution, however, presents a unique challenge rooted in the data-generating process itself. As described in the dataset documentation [4], TWF involves two sequential random checks. Firstly, each machine is assigned a random wear threshold between 200–240 minutes. Secondly, upon crossing that threshold, whether the machine actually fails or receives a tool replacement is determined by another independent random assignment (46 failures out of 120 threshold crossings in the entire dataset). The first random check can in principle be approximated by a probabilistic model. However, the second check is entirely independent of all sensor readings. This introduces irreducible noise which cannot be resolved by any deterministic classification model. As a result, the model achieves lower performance on TWF attribution compared to other failure types. This is expected given the stochasticity in the data-generating process. For the remaining three failure types, which are governed by deterministic threshold rules on observable sensor features, the Bayesian SCM provides accurate attribution.
4.1.2. Generalization and Overfitting Assessment
We performed the multi-seed validation tactic as explained in 3.2.5. Table 8 shows training savings percentage, test savings percentage, and the gap in percentage points for predictive models.
Model Train Savings % Test Savings % Gap (%)
L1: Balanced LogReg 57.2% 51.7% 5.44
L2: Cost-Aware LogReg 56.0% 50.7% 5.26
L3: Cost-Aware Tree 69.4% 65.0% 4.36
L4: Random Forest 73.3% 70.8% 2.44
L5: AdaBoost 69.6% 67.8% 1.82
L6: LightGBM 72.6% 69.0% 3.60
L7: Bayesian SCM 69.5% 66.4% 3.17
Table 8. Train–test savings comparison averaged over validation splits. Gap represents the difference between training and test savings percentages; smaller gaps indicate better generalization.
Most models exhibit modest generalization gaps below 5 percentage points, which we set as an acceptable bound because plant operators typically budget that amount of degradation when deploying predictive maintenance models across production lines [8]. Notably, logistic regression variants (L1, L2) are exceptions which marginally exceed the acceptable bound at 5.44 and 5.26 percentage points respectively. This failure to generalize is not surprising given the inconsistency between their linear decision boundary and the nonlinear ground truth. The remaining models all fall within the acceptable range. AdaBoost (L5) records the smallest gap at 1.82 percentage points, and Random Forest (L4) follows at 2.44 points. The Bayesian SCM (L7) achieves a 3.17 point gap, consistent with the expectation that causal relationships are less susceptible to distributional idiosyncrasies in the training sample. The multi-seed validation confirmed stable model rankings across all five splits, demonstrating that observed performance differences reflect genuine model capabilities rather than chance variation in data partitioning.
5. Conclusion
This study evaluated eight predictive maintenance models on a CNC machine failure dataset, comparing traditional correlation-based machine learning approaches with Bayesian structural causal methods. Among correlation-based models, ensemble methods consistently achieve the best performance, with Random Forest (L4) achieving the best cost reduction of 70.8%. The Bayesian Structural Causal Model (L7) achieved competitive financial performance of a 66.4% cost reduction. On the other hand, it is capable of attributing predicted failures to specific types without requiring failure-type labels during training. This capability can prove to be a favorable trade-off in predictive maintenance scenarios.
This work demonstrates several important findings for predictive maintenance applications:
  1. (1)
    Business-aligned optimization: Models optimized explicitly for business costs (accounting for the 50:1 cost asymmetry between false negatives and false positives) substantially outperform accuracy-optimized approaches.
  2. (2)
    Ensemble methods as performance ceiling: Tree-based ensemble models (L4, L5, L6) consistently outperform single models including the decision tree that mirrors the true data-generating process. This demonstrates the strength of ensembling methods in reducing variance and preventing overfitting.
  3. (3)
    Causal AI for interpretability: The Bayesian SCM (L7) delivers competitive cost savings and provides actionable insights in the form of failure attribution. The performance gap relative to the best ensemble model is modest and may be outweighed by the practical value of the failure diagnosis.
  4. (4)
    Failure attribution without attribution labels: The Bayesian SCM (L7) correctly attributes predicted failures to specific physical mechanisms (TWF, HDF, PWF, OSF) despite being trained only on binary failure labels. This means there is no additional labeling cost required for training the model.
  5. (5)
    Robust generalization: The Bayesian SCM demonstrates solid generalization (3.17 percentage point train-test gap), which is consistent with the expectation that causal relationships are invariant properties of the physical system and less susceptible to distributional variation in training data.
Several limitations of this study should be acknowledged:
  1. (1)
    Dataset scope: The analysis uses a single dataset of 10,000 CNC machines. Generalization to other machine types, manufacturing environments, or operational conditions requires additional validation.
  2. (2)
    Causal graph construction: The causal DAG was manually constructed based on domain knowledge of CNC machine failure mechanisms. In domains where such expertise is unavailable, causal discovery algorithms could be explored, though they introduce additional challenges.
  3. (3)
    Cost estimates: The business costs (25,000 USD for failures, 5,000 USD for preventive maintenance, 500 USD for false alarms) are representative estimates. Actual costs vary by industry, machine criticality, and operational context.
  4. (4)
    Static analysis: This study evaluates models on synthetic data without real-world deployment. Practical implementation would require monitoring for distribution drift, model degradation, and integration with existing maintenance management systems.
  5. (5)
    Class imbalance: The 3.3% failure rate, while realistic for well-maintained equipment, limits the number of positive examples available for model training. Larger datasets with more failure observations could improve model performance.
Future research directions include: (1) validation on operational data; (2) investigation of causal discovery algorithms to automate DAG construction; (3) development of online learning approaches for continual model adaptation; (4) integration with IoT sensor networks and SCADA systems for real-time deployment; and (5) extension to multi-component systems with cascading failure modes.
6. Author Contributions
The authors are listed in alphabetical order.
Krishna Taduri performed exploratory data analysis, developed the causal structural model (including failure attribution) and integrated cost optimization.
Shaunak Dhande performed exploratory data analysis and developed the correlation-based models.
Chutian Ma reviewed and revised the manuscript, and contributed to the discussion of the failure attribution mechanism.
Giacinto Paolo Saggese and Paul Smith provided overall conceptual guidance, contributed to architectural design decisions, and offered feedback on the manuscript.